A 10.64 mol sample of methane gas is maintained in a 0.8073 L container at 301.3 K. What is the pressure in atm calculated using the Van der waal’s equation for CH4 gas under these conditions? For CH4 a = 2.253L ^2 atm/mol^2 and b = 4.278 x 10^-2 L/mol



1
Expert's answer
2021-03-03T02:00:06-0500

Ans:-

Given "V=0.8073 L" "T=301.3 K" "n=10.64 mol" "P=?"

"a= 2.253 \dfrac{L^2 atm}{mol^2}" "b = 4.278\times10^{-2} \dfrac{L}{mol}"


We know that the Vander waal's equation for "CH_4" gas

"P=\dfrac{n\times{R}\times{T}}{V-n\times{b}}-\dfrac{a\times{n^2}}{V^2}"

"\Rightarrow" "P=\dfrac{10.64\times{0.083}\times{301.3}}{0.8073-10.64\times{0.04278}}-\dfrac{2.253\times{10.64^2}}{0.8073^2}"

"\Rightarrow" "P=755.6628-391.3582=364.304atm"

Therefore the pressure of gas "CH_4" is "364.304" atm